How to find variance of a sample


1
Expert's answer
2020-09-17T15:47:36-0400

Let's assume that we have a sample "X_1, X_2, ..., X_n".

The variance of a sample is

"\sigma^2 = Var[X] = \displaystyle \frac{1}{n-1} \sum_{i=1}^n (X_i - \mu)^2"

where "\mu =\displaystyle \frac{1}{n} \sum_{i=1}^n X_i"